HOW DO I CALCULATE?

Saras Salsa Company produces its condiments in two types: Extra Fine for restaurant customers and Family Style for home use.Saras Salsa Company produces its condiments in two types: Extra Fine for restaurant customers and Family Style for home use.
Answer #1
1) Plantwide overhead rate
Estimated overhead/total cases
578250/120,000
4.82
plantwide overhead rate $4.82 per case
total cost per unit
2) Extra Fine 10.82
Family Style 9.82
3) Extra Fine profit 7.18
Family Style loss -0.82
(a) (b) c=a/b Extra fine e=c*d Family style g=c*f
overhead Avtivity Driver Activity rate Activity Drivers overhead Activity Drivers overhead
4) Activity cost (d) assinged (f) assinged
Mixing & cooking 15,100 1,000 machine hours 15.1 per machine hour 300 4530 700 10570
product testing 112,600 600 Batches 187.6667 per batch 210 39410 390 73190
Machine Calibration 255,000 500 production run 510 per production run 210 107100 290 147900
Labeling & Defects 19,500 130,000 cases of output 0.15 per case 21000 3150 109,000 16350
Recipe formulation 91,000 50 Focus groups 1820 per focus group 32 58240 18 32760
heat,light and water 28,000 1,000 machine hours 28 per machine hour 300 8400 700 19600
Material handling 66,000 8 container type 8250 per container type 7 57750 1 8250
587,200 total OH cost assign 278580 total OH cost assign 308620
total units produced 21,000 total units produced 109,000
oh cost per unit 13.27 oh cost per unit 2.83
Extra Family
fine Style
Direct materials & direct labor 4.00 3.00
overhead 13.27 2.83
tota cost/case 17.27 5.83
5) Extra Fine loss -2.27 (15-17.27)
Family Style profit 0.17 (6-5.83)